Pure tin was mixed with an 8% tin alloy to produce an alloy that was 25% tin. How much pure tin and how much 8% alloy were used to produce 184 Kg of 25%alloy.
Answer by mananth(16946)

percent ---------------- quantity
Pure Tin 100.00% ---------------- x kg
Tin impure 8.00% ------ 184 - x kg
Mixture 25.00% ---------------- 184
12
100.00% x+8.00%( 184-x)= 25.00%* 184
100x+8(184-x)=4600
100x+1472-8x=4600
100x-8x= 4600 - (-1472 )
92x=3128
/ 92
x=34kg 100.00% Pure Tin
150kg 8.00% Tin impure

m.ananth@hotmail.ca